OpenStudy (anonymous):

Help plzzz: In a triangle PQR, 2 of the sides are 5 units long and the included angle is 100°. How long is the third side?

OpenStudy (anonymous):



OpenStudy (anonymous):

you know cosine formula??

OpenStudy (anonymous):

yea i do but i dont understand it very well

OpenStudy (anonymous):

see it's like, \[a^ 2 = b ^2 + c ^2 − 2bc \cos A\] where both sides b and c has been given and angle A is also given, just need to substitute in it

OpenStudy (anonymous):

okayy ill try it out

OpenStudy (anonymous):

yeah!! go on )

OpenStudy (anonymous):

So is this right?\[5^2 = 5^2 + c^2 - 2 * 5c * Cos(100)\]

OpenStudy (anonymous):

when substituted?

OpenStudy (anonymous):

oh!!! no actually from my equation you need to calculate "a"...your b and c are the terms are given )

OpenStudy (anonymous):

oh yea course so \[a^2 = 5^2+5^2-2*5*5*\cos(100)\]

OpenStudy (anonymous):

yeah that one @FishFingersandCustard Now calculate a from that equation )

OpenStudy (anonymous):

thanks so much @aryandecoolest helped so much :D

OpenStudy (anonymous):

no problem ) pleasure to help :D @FishFingersandCustard